### 13 Mar 2026 v1

@@ Abstract
Early Pruning can be integrated with minimal changes to existing codebases and requires only a one-time preprocessing step. Across multiple state-of-the-art RAPTOR-based solutions, including RAPTOR, ULTRA-RAPTOR, McRAPTOR, BM-RAPTOR, ULTRA-McRAPTOR, and UBM-RAPTOR and tested on the Switzerland and London transit networks, we achieved query time reductions of up to 57%. This approach provides a generalizable improvement to the efficiency of transit pathfinding algorithms.
@@ 2 Literature Review
The development of efficient algorithms for public transit routing has been a sustained research effort spanning over two decades. Early approaches adapted Dijkstra’s algorithm Dijkstra [1959] and its extensions to the transit network Brodal and Jacob [2003], Rohovyi et al. [2024, 2025] for time-dependent or time-expanded graph models, but the resulting representations were often large and inefficient for the temporal structure of timetable data. These limitations motivated algorithms that operate directly on timetable data, including the Connection Scan Algorithm (CSA) Dibbelt et al. [2017], Trip-Based Routing Köhler et al. [2015], Transfer Patterns Bast et al. [2010], and TCD Abuaisha et al. [2024]. Among these, RAPTOR Delling et al. [2012] has achieved the broadest practical adoption due to its round-based design, which naturally produces Pareto-optimal journeys. The RAPTOR framework has since been extended in many directions: MR and MCR Delling et al. [2013] enable unlimited transfers via Dijkstra-based transfer relaxation; BM-RAPTOR Delling et al. [2019] and UBM-RAPTOR Sauer [2024] constrain the Pareto space for faster queries with additional criteria; and the ULTRA framework Baum et al. [2019] precomputes shortcut edges to avoid transitive closure entirely. Wagner and Zündorf Wagner and Zündorf [2017] demonstrated the significant impact of unrestricted walking on travel times using a profile algorithm based on MCR. Phan and Viennot Phan and Viennot [2019] subsequently explored hub labeling as an alternative approach, decomposing walking transfers into two consecutive hops and achieving significant speedups for both RAPTOR and CSA-based algorithms.
@@
Pruning strategies are well-established in the transit routing literature. The original RAPTOR paper Delling et al. [2012] introduced several forms of pruning: marking (only processing routes that serve recently updated stops), local pruning (discarding labels dominated at a stop), and target pruning (discarding labels dominated by the best known arrival at the target). Our Early Pruning technique is most closely related to target pruning, but operates at a different granularity. While standard target pruning checks whether a label at a stop is dominated by the target’s best known arrival, Early Pruning exploits the ordering of transfer edges to terminate the entire transfer iteration at a stop early. This is possible because pre-sorting transfers by duration guarantees that all subsequent transfers will produce later arrivals than the current one. The same principle was introduced by Phan and Viennot [2019] in the context of hub labellings, where sorted hub lists enable a target-pruning optimisation that stops scanning once arrival times exceed the best-known target time. Our work applies this principle directly within the transfer relaxation loop of RAPTOR, without requiring hub label preprocessing, and extends it to settings with, additional criteria, such as costs or walking duration, including McRAPTOR, BM-RAPTOR, and their ULTRA-based counterparts.
@@ 3.2 Algorithms
McRAPTOR Delling et al. [2012] extends RAPTOR to optimise additional criteria beyond arrival time and number of transfers, such as walking duration and costs. Given a set of criteria, a journey $J$ dominates another journey $J^{\prime}$, if $J$ is not worse than $J^{\prime}$ in any criterion. In our research, we use a three-criteria version of McRAPTOR, which includes walking duration, arrival time, and the number of transfers. Although McRAPTOR can be very useful in practice, as it offers users various options that can be sorted and selected according to their preferences, it has a long query time. One way to speed it up is to find the solution not within the entire search space, but within a defined, restricted Pareto space. This modification, called BM-RAPTOR Delling et al. [2019], reduces the query time by 65 times compared to McRAPTOR, which is still roughly twice as slow as standard RAPTOR Sauer [2024].
@@ 4 Early Pruning
The same idea can be applied to the extended-criteria versions of the algorithm, such as McRAPTOR, ULTRA-McRAPTOR, BM-RAPTOR, and UBM-RAPTOR. However, in this case, we use dominance instead of direct scalar comparison. Since our research employs three criteria—arrival time, walking duration, and number of transfers—the goal of Early Pruning is to skip transfers that cannot possibly yield a non-dominated result. Because the transfers are pre-sorted by duration, if a new partial path (e.g., arrival at stop $b$) is already dominated by an existing path to target $t$ in the Pareto sense (meaning the new path is worse in all criteria), the increasing walking duration makes it impossible to dominate in any of the listed criteria, so we can safely prune that path and all subsequent, longer transfers.
@@ 5 Experiments
The experiments were conducted on the London and Switzerland networks, which were previously accessible from the Karlsruhe Institute of Technology (KIT). [^5] https://i11www.iti.kit.edu/PublicTransitData/ULTRA/. KIT Baum et al. [2019] constructed Unrestricted Transfer Graphs by extracting road graphs, including pedestrian zones and staircases, from OpenStreetMap [^6] https://download.geofabrik.de/. Unless stated otherwise, they used walking as the transfer mode, assuming a constant speed of 4.5 km/h. However, the Unrestricted Transfer Graph cannot be used with algorithms that require a transitively closed graph, since the transitive closure would be too large Wagner and Zündorf [2017]. To address this, the authors precomputed more compact Transitive Transfer Graphs by adding edges between all stops whose transfer-graph distance was below a chosen threshold (9 minutes for Switzerland and 4 minutes for London) and then computing the transitive closure. The thresholds were set so that the resulting graph had an average vertex degree of about 100 Baum et al. [2019]. Finally, two types of ULTRA precomputations were performed on the Unrestricted Transfer Graphs to generate shortcuts: one for earliest-arrival algorithms and another for multicriteria algorithms. Networks statistics used in the experiments are provided in Table 1.
@@
The runtime reductions we report translate directly into operational benefits for real-world journey planners. For a metropolitan public-transport system that handles millions of trip queries per day, a 10–30% decrease in average query time allows agencies to support more complex multimodal options—such as flexible walking transfers or shared-mobility connections—without increasing server capacity. This improves the responsiveness of passenger-facing apps, enables planners to expand transfer radii to capture more realistic travel choices, and lowers infrastructure and energy costs for operators. Algorithms that do not require additional preprocessing, unlike ULTRA-based methods, benefit more from Early Pruning. Consequently, this technique is especially effective for dynamic algorithms operating in changing environments.
@@ 6 Implications for Transport Policy and Planning
Incorporating additional transfer modes into journey planners, as envisioned by Mobility-as-a-Service (MaaS) platforms that integrate walking, cycling, e-scooters, and shared mobility into a single interface, increases the density of the transfer graph, which directly increases query times. In practice, transit agencies must choose between offering comprehensive multimodal options and maintaining fast query responses. Early Pruning relaxes this trade-off: the 34% speedup achieved for RAPTOR on the London transitive graph could allow an agency to increase the distance threshold used to compute the transitive closure from 4 minutes to a longer duration while maintaining the same query time, thereby including a richer set of transfer connections between stops. For queries that optimise additional criteria beyond arrival time and number of transfers, such as walking duration or costs, the 57% speedup achieved for McRAPTOR means that a planner returning results in 4 seconds can now respond in under 2 seconds, bringing it within the range of interactive applications.
@@ 7 Conclusion and Future Work
In this work, we introduce a new technique, Early Pruning, which accelerates various pathfinding algorithms in public transit routing. We adapted it for several RAPTOR modifications and observed consistent improvements across all variants, with speedups ranging from 2% to 57%, depending on network characteristics and algorithm modifications. In particular, denser networks benefit more from this approach. By enabling a speedup of up to 57% on highly complex queries (like McRAPTOR), which provide users with multiple multimodal transfer options, Early Pruning makes advanced multimodal, multi-criteria planning computationally feasible for large-scale production systems like trip planners. This not only results in a better user experience but can also potentially influence mode choice by delivering richer, faster routing information.

### 21 Apr 2026 v2

@@ Abstract
Early Pruning can be integrated with minimal changes to existing codebases and requires only a one-time preprocessing step. The technique preserves Pareto-optimality in extended-criteria settings whenever the additional optimization criteria are monotonically non-decreasing in transfer duration. Across multiple state-of-the-art RAPTOR-based solutions, including RAPTOR, ULTRA-RAPTOR, McRAPTOR, BM-RAPTOR, ULTRA-McRAPTOR, and UBM-RAPTOR and tested on the Switzerland and London transit networks, we achieved query time reductions of up to 57%. This approach provides a generalizable improvement to the efficiency of transit pathfinding algorithms.
@@ 2 Literature Review
The development of efficient algorithms for public transit routing has been a sustained research effort spanning over two decades. Early approaches adapted Dijkstra’s algorithm Dijkstra [1959] and its extensions to the transit network Brodal and Jacob [2003], Rohovyi et al. [2024] for time-dependent or time-expanded graph models, but the resulting representations were often large and inefficient for the temporal structure of timetable data. These limitations motivated algorithms that operate directly on timetable data, including the Connection Scan Algorithm (CSA) Dibbelt et al. [2017], Trip-Based Routing Witt [2015, 2016], Transfer Patterns Bast et al. [2010], and TCD Abuaisha et al. [2024]. Among these, RAPTOR Delling et al. [2012] has achieved the broadest practical adoption due to its round-based design, which naturally produces Pareto-optimal journeys. The RAPTOR framework has since been extended in many directions: MR-$\infty$ (MR) and MCR Delling et al. [2013] enable unlimited transfers via Dijkstra-based transfer relaxation; BM-RAPTOR Delling et al. [2019] and UBM-RAPTOR Sauer [2024] constrain the Pareto space for faster queries with additional criteria; and the ULTRA framework Baum et al. [2019] precomputes shortcut edges to avoid transitive closure entirely. Wagner and Zündorf Wagner and Zündorf [2017] demonstrated the significant impact of unrestricted walking on travel times using a profile algorithm based on MCR. Phan and Viennot Phan and Viennot [2019] subsequently explored hub labeling as an alternative approach, decomposing walking transfers into two consecutive hops and achieving significant speedups for both RAPTOR and CSA-based algorithms. Dijkstra-based approaches have also continued to evolve in parallel, with recent work addressing unlimited-transfer routing Rohovyi et al. [2025], Katkalo et al. [2026]
@@
Pruning strategies are well-established in the transit routing literature. The original RAPTOR paper Delling et al. [2012] introduced several forms of pruning: marking (only processing routes that serve recently updated stops), local pruning (discarding labels dominated at a stop), and target pruning (discarding labels dominated by the best known arrival at the target). Our Early Pruning technique is most closely related to target pruning, but operates at a different granularity. While standard target pruning checks whether a label at a stop is dominated by the target’s best known arrival, Early Pruning exploits the ordering of transfer edges to terminate the entire transfer iteration at a stop early. This is possible because pre-sorting transfers by duration guarantees that all subsequent transfers will produce later arrivals than the current one. The same principle was introduced by Phan and Viennot [2019] in the context of hub labellings, where sorted hub lists enable a target-pruning optimisation that stops scanning once arrival times exceed the best-known target time. Our work applies this principle directly within the transfer relaxation loop of RAPTOR, without requiring hub label preprocessing, and extends it to settings with additional criteria, such as walking duration, including McRAPTOR, BM-RAPTOR, and their ULTRA-based counterparts.
@@ 3.2 Algorithms
McRAPTOR Delling et al. [2012] extends RAPTOR to optimise additional criteria beyond arrival time and number of transfers, such as walking duration. Given a set of criteria, a journey $J$ dominates another journey $J^{\prime}$, if $J$ is not worse than $J^{\prime}$ in any criterion. In our research, we use a three-criteria version of McRAPTOR, which includes walking duration, arrival time, and the number of transfers. Although McRAPTOR can be very useful in practice, as it offers users various options that can be sorted and selected according to their preferences, it has a long query time. One way to speed it up is to find the solution not within the entire search space, but within a defined, restricted Pareto space. This modification, called BM-RAPTOR Delling et al. [2019], reduces the query time by 65 times compared to McRAPTOR, which is still roughly twice as slow as standard RAPTOR Sauer [2024].
@@ 4 Early Pruning
The same idea extends to extended-criteria variants such as McRAPTOR, ULTRA-McRAPTOR, BM-RAPTOR, and UBM-RAPTOR, provided a monotonicity condition holds: each additional criterion beyond arrival time and number of transfers must be monotonically non-decreasing in the transfer duration. Under this condition, once a partial path from stop $u$ is dominated in the Pareto sense by an existing solution to target $t$, all subsequent transfers (which are longer, due to pre-sorting) will produce paths that are at least as bad in every criterion. Thus they can be safely pruned.
Remark 1. Let $s$ be a stop with arrival time $\tau_{k}(s)$, and let $e_{i}=(s,v_{i})$ and $e_{j}=(s,v_{j})$ be two outgoing transfers with $\tau(e_{i})\leq\tau(e_{j})$. Early Pruning preserves Pareto-optimality if for every optimization criterion $c\in C$, the label produced by taking transfer $e_{j}$ is not better than the label produced by taking transfer $e_{i}$. That is, a longer transfer from the same stop cannot improve any criterion.
In our experiments, we use three criteria: arrival time, walking duration, and number of transfers. A longer transfer from the same stop leads to a later arrival time and a longer walk, while the number of transfers remains unchanged. All three criteria therefore satisfy the monotonicity condition from Remark 1. The same holds for other transfer mode durations, such as cycling time or e-scooter travel time, making Early Pruning applicable to extended-criteria formulations beyond the specific configuration tested here.
@@ 5 Experiments
The experiments were conducted on the London and Switzerland networks, which were previously accessible from the Karlsruhe Institute of Technology (KIT). [^5] https://i11www.iti.kit.edu/PublicTransitData/ULTRA/. KIT Baum et al. [2019] constructed Unrestricted Transfer Graphs by extracting road graphs, including pedestrian zones and staircases, from OpenStreetMap [^6] https://download.geofabrik.de/. Unless stated otherwise, they used walking as the transfer mode, assuming a constant speed of 4.5 km/h. However, the Unrestricted Transfer Graph cannot be used with algorithms that require a transitively closed graph, since the transitive closure would be too large Wagner and Zündorf [2017]. To address this, the authors precomputed more compact Transitive Transfer Graphs by adding edges between all stops whose transfer-graph distance was below a chosen threshold (9 minutes for Switzerland and 4 minutes for London) and then computing the transitive closure. The thresholds were set so that the resulting graph had an average vertex degree of about 100 Baum et al. [2019]. Finally, two types of ULTRA precomputations were performed on the Unrestricted Transfer Graphs to generate shortcuts: one for earliest-arrival algorithms and another for extended-criteria algorithms. Networks statistics used in the experiments are provided in Table 1.
@@
The runtime reductions we report translate directly into operational benefits for real-world journey planners. For a metropolitan public-transport system that handles millions of trip queries per day, a 10–30% decrease in average query time allows agencies to support more complex multimodal options, such as flexible walking transfers or shared-mobility connections, without increasing server capacity. This improves the responsiveness of passenger-facing apps, enables planners to expand transfer radii to capture more realistic travel choices, and lowers infrastructure and energy costs for operators. Algorithms that do not require additional preprocessing, unlike ULTRA-based methods, benefit more from Early Pruning. Consequently, this technique is especially effective for dynamic algorithms operating in changing environments.
@@ 6 Implications for Transport Policy and Planning
Incorporating additional transfer modes into journey planners, as envisioned by Mobility-as-a-Service (MaaS) platforms that integrate walking, cycling, e-scooters, and shared mobility into a single interface, increases the density of the transfer graph, which directly increases query times. In practice, transit agencies must choose between offering comprehensive multimodal options and maintaining fast query responses. Early Pruning relaxes this trade-off: the 34% speedup achieved for RAPTOR on the London transitive graph could allow an agency to increase the distance threshold used to compute the transitive closure from 4 minutes to a longer duration while maintaining the same query time, thereby including a richer set of transfer connections between stops. For queries that optimise additional criteria beyond arrival time and number of transfers, such as walking duration, the 57% speedup achieved for McRAPTOR means that a planner returning results in 4 seconds can now respond in under 2 seconds, bringing it within the range of interactive applications.
@@ 7 Conclusion and Future Work
In this work, we introduce a new technique, Early Pruning, which accelerates various pathfinding algorithms in public transit routing. We adapted it for several RAPTOR modifications and observed consistent improvements across all variants, with speedups ranging from 2% to 57%, depending on network characteristics and algorithm modifications. In particular, denser networks benefit more from this approach. By enabling a speedup of up to 57% on highly complex queries (like McRAPTOR), which provide users with additional criteria such as walking duration, Early Pruning makes advanced multimodal planning computationally feasible for large-scale production systems like trip planners. This not only results in a better user experience but can also potentially influence mode choice by delivering richer, faster routing information.